An airplane flew for 4hr with a 20-mph tailwind. The return flight against the same wind took 5 hr. Find the speed of the plane in still air.
Can anyone help me? I do not know how to even start.

Found 2 solutions by sridhar, stanbon:
Answer by sridhar(2)


S = V * T
S = (V + 20) * 4
S = (v - 20) * 5
So, (V + 20) * 4 = (v - 20) * 5
4V + 80 = 5V - 100
v = 180
Speed of the plane = 180 mph

Answer by stanbon(75887)

An airplane flew for 4hr with a 20-mph tailwind. The return flight against the same wind took 5 hr. Find the speed of the plane in still air.
----------------------------
With the wind DATA:
time = 4 hr: distance = x miles; rate = d/t = x/4 mph
-------------------------------
Against the wind DATA:
time = 5 hr. distance = x miles ; rate = x/5 mph
-------------------------------------------------------
Equations:
p + 20 = x/4
p - 20 = x/5
-----------------
2p = (x/4)+(x/5)
2p = 9x/20
p = 9x/40
------------
Substitute into p + 20 = x/4 to solve for "x"
(9x/40) + 20 = (x/4)
Multiply thru by 40 to get:
9x + 800 = 10x
x = 800
x = 800 (distance traveled)
-----------------
Subtitute to solve for "p":
p = x/4 = 800/4 = 200 mph (plane's speed in still air)
=======================================
Cheers
Stan H.